Three capacitors of capacitances C1 = 15 μF, C2 = 30 μF, and C3 = 45μF were connected in series to a potential difference V. Find the equivalent capacitance of the three capacitors.

1
Expert's answer
2022-04-15T16:18:03-0400

"\frac 1C=\frac 1{C_1}+\frac 1 {C_2}+\frac 1{C_3},"

"C=8.2~\mu F."